Compute the divergence of the vector field exy2i+x+2yj



Select one:



A. exy2+2




B. y2+2




C. 2exy




D. exy2+2x+y


1
Expert's answer
2022-04-04T16:13:09-0400

"F=e^xy^2i+\left( x+2y \right) j\\divF=\frac{\partial \left( e^xy^2 \right)}{\partial x}+\frac{\partial \left( x+2y \right)}{\partial y}=e^xy^2+2\\A\,\,is\,\,correct"